OpenStudy (anonymous):

solve for b L=b-c/2 I need to check my answer to make sure it's correct TY!

OpenStudy (anonymous):

is it 2L+c

OpenStudy (anonymous):

No it is not.

OpenStudy (anonymous):

:(

OpenStudy (oaktree):

No it totally is 2L+c.

OpenStudy (anonymous):

\[L = b - \frac{c}{2}\]\[L + \frac{c}{2} = b\] That's all you need to do. You can't clear the fraction because you have to multiply both sides by 2, which would result in \(2L + c = 2b\) which is no longer solved for b.

OpenStudy (oaktree):

No the entire expression (b-c) is being divided by 2 I think. @mallory2023, confirm or deny?

OpenStudy (anonymous):

you are correct oaktree ty, I second guessed myself

OpenStudy (anonymous):

Ok. If it's L = (b-c)/2 then it is correct. \[L = \frac{b-c}{2} \\ 2L = b -c \\ 2L + c = b\]As written above.

OpenStudy (anonymous):